# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ 9 Concluding remarks
In 1971 Migdal suggested possibility of existence of metastable or stable abnormal pion condensate nuclei and Bodmer assumed possibility of collapsed quark nuclei. In 1974 Lee and Wick proposed scalar (sigma) condensate abnormal nuclei. In 1984 Witten suggested quark nuggets. These pioneering suggestions paved the way for numerous subsequent theoretical and experimental investigations of various anomalous nuclear objects. New more detailed models were constructed. Nevertheless, in spite of more than fifty years were past, as in early days of this field, there remain essential uncertainties in theoretical predictions due to a lack of our knowledge of a strongly interacting dense matter. In spite of abnormal nuclear states were not observed, the number of observations grows, which are hardly explained by conventional physics. Some of such events could be associated with manifestation of anomalous nuclear systems. Let us be patient. Time flows not in vain.
[…]
All authors have equally contributed to the research and preparation of the manuscript.
[…]
@@
EEK acknowledges the support by the Slovak grant VEGA 1/0521/22 and the support within the framework of the scientific program of the National Center for Physics and Mathematics in Sarov, topic No. 6 "Nuclear and Radiation Physics" (stage 2023-2025).
[…]
Not applicable
[…]
Not applicable
[…]
No explicit data set are prepared.
[…]
The authors declare no conflict of interest.
[…]
Abbreviations The following abbreviations are used in this manuscript:
EoS Equation of State
HESS High Energy Stereoscopic System
ISM Isospin-Symmetric Matter
NICER Neutron Star Interior Composition Explorer
PCAC Partial Conservation of Axial Current
PNM Pure Neutron Matter
PSR Pulsating Source of Radiation
QCD Quantum Chromo-Dynamics
RMF Relativistic Mean Field
XTE Rossi X-Ray Timing Explorer
4U Fourth UHURU Catalogue
[…]

### v2

@@ IX Conclusion
Concluding, in spite of more than fifty years were past, as in early days of this field, there remain essential uncertainties in theoretical predictions due to a lack of our knowledge of the behavior of the strongly interacting dense and dilute nuclear matter. In spite of abnormal nuclear states were not observed, the number of observations grows, which are hardly explained by conventional physics. Some of such events could be associated with manifestation of anomalous nuclear systems. Let us be patient. Time flows not in vain.
Acknowledgment
@@
EEK acknowledges the support by the Slovak grant VEGA 1/0521/22 and the support within the framework of the scientific program of the National Center for Physics and Mathematics in Sarov, topic No. 6 ”Nuclear and Radiation Physics” (stage 2023-2025).